OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6)

OpenStudy (anonymous):

I get -2x + -2y = 26 and that is not one of my choices

OpenStudy (decentnabeel):

first find the slope.. do you find the slope

OpenStudy (anonymous):

0.3?

OpenStudy (anonymous):

Im sorry but im getting -2,-2

OpenStudy (anonymous):

so would the slope be 0?

OpenStudy (decentnabeel):

slope=y2-y1/(x2-x1) the point is the given (-5,-7)(-8,-6) x1=-5 y1=-7 x2=-8 y2=-6 plug in .. slope=m=(-6+7)/(-8+5) m=-1/3

OpenStudy (anonymous):

I apologize but im still not understanding

OpenStudy (decentnabeel):

do you point slope formula..

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (decentnabeel):

slope=m m=(y2-y1)/(x2-x1)

OpenStudy (decentnabeel):

ok.. first find the slope

OpenStudy (anonymous):

-1/3

OpenStudy (decentnabeel):

m=-1/3 so use the point (-5,-7) x1=-5 y1=-7 y - y1 = m(x - x1 y+7=(-1/3)(x+5) \[y+7=-\frac{1}{3}x-\frac{5}{3}\]

OpenStudy (decentnabeel):

yes good slope =-1/3

OpenStudy (anonymous):

y + 7 = -1/3 x - 5 or y + 6 = -1/3 x + 7 this is the answer I got after plugging everything in

OpenStudy (decentnabeel):

y + 7 = -1/3 x - 5/3 not only 5

OpenStudy (decentnabeel):

\[y+6=-\frac{1}{3}\left(x+5\right)\] -1/3 multiple for both x and 5

OpenStudy (anonymous):

I got it thnk you for all your help